OpenStudy (anonymous):

Carl can paint a house in 8 hours. After he had been working for 1 hour, His brother Gomer started helping him. It then took them 3 hours to finish. How long would it take Gomer to paint a house by himself?

OpenStudy (anonymous):

Look at rates of work. Let's say that Gomer can paint a house in x hours. Then his rate is 1/x house per hour. Carl works at a rate of 1/8 house per hour. They worked together on the house for 2 hours. set up an expression representing the fraction of the house they did during during the two hours they worked together. Add to this the fraction Carl did when working alone. This is the fraction that was done during the 3 hours. Put this expression equal to 1 because they finished the entire house.

OpenStudy (anonymous):

So the equation would be 1/8 + 1/x = 1/3 ??

OpenStudy (anonymous):

The equation is 1/8+1/8(2)+(1/x)(2)=1 which gives you 2/x=5/8